OpenStudy (anonymous):

Write an equation of the line that has a slope of -3 and contains the point (1, 1) in point-slope form.

OpenStudy (anonymous):

(y1-y2)=m(x1-x2)....Where m = slope and (x2,y2) are the points on the line.